OpenStudy (anonymous):

Find an expression for the function whose graph is the bottom half of the parabola x + (y - 8)2 = 0.

OpenStudy (anonymous):

is it -sqrt(x-8)^2

OpenStudy (watchmath):

\((y-8)^2=-x\) \(y-8=\pm\sqrt{-x}\) For the bottom half, \(y<8\) So \(y=8- \sqrt{-x}\)